. Travel Co. specializes in trans-American moves for businesses and individuals through the shipment of containers. For every container the company transports via rail, Travel Co. must pay the rail owners $3,000. In addition, containers are weighed and for each pound of weight, the company must pay $0.20. Harry plans to charge $0.60 per pound to his customers.
How many pounds would a container have to hold in order for the company to Break Even? (10 pts)
What would the total revenue be at Break Even? (5 pts)
The company finds out that to maintain the integrity of its containers, each container could only hold an average of 15,000 pounds. What would the company have to charge customers to make $5,000 in profit on each container? (10 pts)
Answer #1

1st part,

Fixed cost = $3000

Variable cost = $0.20/pound

Revenue = $0.60/pound

How many pounds would a container have to hold in order for the company to Break Even?

Let, required quantity = x pound

To break-even,
3000+0.2x = 0.6x

or, x = 3000/0.4 = 7500

What would the total revenue be at Break Even?

total revenue be at Break Eve = 7500*0.6 = 4500

What would the company have to charge customers to make $5,000 in profit on each container?

Total cost = 3000+15000*0.2 = 6000

Let, Required selling price = x/pound

6000+5000 = x*15000

or, x = 11000/15000 = $0.733333333/pound